Grade 12Algebra

Check whether the value given in the brackets is a solution to the given equation or not:
4p – 3 = 13 (p = 1)

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

LHS = 4p - 3
By substituting the value of p = 1
Then,
LHS = 4p - 3
= (4 × 1) - 3
= 4 - 3
=1
By comparing LHS and RHS 1≠13
LHS≠RHS
Hence, the value of p = 1 is not a solution to the given equation 4p - 3 = 13.

Thanks